A bucket contains 10 balls.  3 are red, 4 are blue and 3 are green?  If you randomly choose 5 balls without replacing any of them, what is that probability that at least one is green?   




1
Expert's answer
2021-05-31T19:08:33-0400
"\dbinom{3}{1}\dbinom{4}{4}=3(1)=3"

"\dbinom{3}{2}\dbinom{4}{3}=3(4)=12"

"\dbinom{3}{3}\dbinom{4}{2}=1(6)=6"

"3+4+3=10"

"\dbinom{10}{5}=252"

"P(at\ least\ 1 \ green)=1-\dfrac{3+12+6}{252}=\dfrac{11}{12}"